OpenStudy (anonymous):

What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (ccswims):

well think of it this way, they are doing the communicative property of addition

OpenStudy (ccswims):

that means such as 12+3+4 = 4+12+3

OpenStudy (ccswims):

so if 12 + (-7) + (-14) what would (-7 + s) + (-14) be?

OpenStudy (ccswims):

what is missing from equation two that equation one has?

OpenStudy (igreen):

(–7 + s) + (–14) = (12 + (–7)) + (–14) Well we can simplify it first. -7 + s - 14 = 12 - 7 - 14 Simplify all whole numbers: s - 21 = -9 If we put in 12 we get: 12 - 21 = -9 -9 = -9 So 12 is your answer.

OpenStudy (ccswims):

there is a much simpler way

OpenStudy (ccswims):

you could've just seen 12 was the only other thing missing

OpenStudy (ccswims):

but which ever way you do it both are correct :)